OpenStudy (anonymous):

differentiate the following equation sqrt(5)/(x^4)

OpenStudy (abb0t):

\(\frac{ \sqrt{5} }{ x^4 } = \sqrt{5} \frac{ d }{ dx }(x^{-4})\) the \(\sqrt{5}\) is a constant and can just be factored out.

OpenStudy (anonymous):

so would my final answer be 4x^3sqrt(5)

OpenStudy (inkyvoyd):

no... follow the power rule for differentiation: \(\frac{d}{dx}x^n=nx^{n-1}\) remember that n=-4

OpenStudy (anonymous):

-4sqrt(5)/x^5

OpenStudy (anonymous):

oh ok can we leave the exponent as a negative

OpenStudy (abb0t):

yes.

OpenStudy (abb0t):

Or you can simplify it to make it look "nicer". Since you should recognize that: \(\huge a^{-n} = \frac{ 1 }{ a^n }\)

OpenStudy (anonymous):

Thanks